A chemist needs 500ml of a solution that is 25% acid. She combines x ml of a 15% acid solution with y ml of a 40% acid solution. Write and solve a system of equations to find the values of x and y.
Answer by Boreal(15235)

x+y=500
.15x+.40y=500(.25), the equation of "pure acid"
multiply the bottom by 100 to clear decimals
15x+40y=12500; multiply the first equation by minus 15 to eliminate x when they are added together.
-15x-15y=-7500
25y=5000
y=200 ml of 40%
x=300 ml of 15%
.15(300)=45
.4(200)=80
That equals 125, which is 500*0.25